# Neural networks for the approximation of Euler’s elastica

### v1

@@ Abstract
Euler’s elastica is a classical model of flexible slender structures, relevant in many industrial applications. Static equilibrium equations can be derived via a variational principle. The accurate approximation of solutions of this problem can be challenging due to nonlinearity and constraints. We here present two neural network based approaches for the simulation of this Euler’s elastica. Starting from a data set of solutions of the discretised static equilibria, we train the neural networks to produce solutions for unseen boundary conditions. We present a discrete approach learning discrete solutions from the discrete data. We then consider a continuous approach using the same training data set, but learning continuous solutions to the problem. We present numerical evidence that the proposed neural networks can effectively approximate configurations of the planar Euler’s elastica for a range of different boundary conditions.
@@ 1 Introduction
Modelling of mechanical systems is relevant in various branches of engineering. Typically, it leads to the formulation of variational problems and differential equations, whose solutions are approximated with numerical techniques. The efficient solution of linear and non-linear systems resulting from the discretisation of mechanical problems has been a persistent challenge of applied mathematics. While classical solvers are characterised by a well-established and mature body of literature [1, 2, 3, 4, 5, 6, 7], the past decade has witnessed a surge in the use of novel machine learning-assisted techniques [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. These approaches aim at enhancing solution methods by leveraging the wealth of available data and known physical principles. The use of deep learning techniques to improve the performance of traditional numerical algorithms in terms of efficiency, accuracy, and computational scalability, is becoming increasingly popular also in computational mechanics. Examples comprise virtually any problem where approximation of functions is required, but also efficient reduced order modelling e.g. in fluid mechanics, the deep Ritz method, or more specific numerical tasks such as optimisation of the quadrature rule for the computation of the finite element stiffness matrix, acceleration of simulations on coarser meshes by learning appropriate collocation points, and replacing expensive numerical computations with data-driven predictions [9, 25, 26, 27, 12]. This recent literature is evidence that neural networks can be used successfully as surrogate models for the solution operators of various differential equations.
@@
In the context of ordinary and partial differential equations, two main trends can be identified. The first one aims at providing a machine learning based approximation to the discrete solutions of differential problems on a certain space-time grid, for example by solving linear or nonlinear systems efficiently and accelerating convergence of iterative schemes [20, 19, 15, 14, 13]. The second one provides instead solutions to the differential problem as continuous (and differentiable) functions of the temporal and spatial variables. Depending on the context, conditions on such approximate solutions are then provided by the differential problem itself, by the initial values and the boundary conditions, and by the available data. The idea of providing approximate solutions as functions defined on the space-time domain and parametrised as neural networks was proposed in the nineties [28] and was recently revived in the framework of Physics-Informed Neural Networks in [10]. Since then, such an approach has attracted a lot of interest and has developed in many directions [8, 21, 25].
@@
In this work, we use neural networks to approximate the configurations of highly flexible slender structures modelled as beams. Such models are of great interest in industrial applications like cable car ropes, diverse types of wires or endoscopes [29, 30, 31, 32]. Notwithstanding their ingenious and simple mathematical formulation, slender structure models can accurately reproduce complex mechanical behaviour and for this reason their numerical discretisation is often challenging. Furthermore, the use of 3-dimensional models requires high computational time. Due to the fact that slender deformable structures have one dimension (length) being orders of magnitude larger than their other dimensions (cross-section), it is possible to reduce the complexity of the problem from a $3$-dimensional elastic continuum to a $1$-dimensional beam. A beam is modelled as a centerline curve, $\mathbf{q}:[0,L]\to\mathbb{R}^{n},s\mapsto\mathbf{q}(s)$, with $n=2$ or $n=3$, along which a rigid cross-section $\Sigma(s)$ is attached. The main model assumption is that the diameter of $\Sigma(s)$ is small compared with the undeformed length $L$. We here consider a special case of a beam where the cross-section $\Sigma(s)$ is constant and orthogonal to the centerline, the $2$-dimensional Euler’s elastica [33]. In this case, $\mathbf{q}(s)$ is inextensible with fixed boundary conditions and is the solution of a bending energy minimisation problem [34, 35, 36].
@@
When approximating static equilibria of the Euler’s elastica via neural networks, a key issue is to ensure the inextensibility of the curve (having unit norm tangents) as well as the boundary conditions. Two main approaches can be found in the literature [25, 21, 37]. One is the weak imposition of constraints and boundary conditions adding appropriate, extra terms to the loss function. The other is a strong imposition strategy consisting in shaping the network architectures so that they satisfy the constraints by construction. We show examples of both the approaches in Sections 4 and 5.
@@
The paper is organised as follows. In Section 2, we present the mathematical model of the planar Euler’s elastica, including its continuous and discrete equilibrium equations. We describe the approach used to generate the data sets for the numerical experiments. In Section 3, we introduce some basic theory and notation for neural networks that we shall use in the succeeding sections. Starting from general theory, we specialise in the task of approximating configurations of the Euler’s elastica. In Section 4, we introduce the discrete approach, which aims to approximate precomputed numerical discretisations of the Euler’s elastica. We discuss some drawbacks associated with this approach and then propose an alternative approximation strategy in Section 5. The continuous approach consists in computing an arc length parametrisation of the beam configuration. We provide insights into two additional networks and analyse how the test accuracy changes with varying constraints, such as boundary conditions or tangent vector norms. Data and codes for the numerical experiments are available in the GitHub repository associated to the paper  [^1] https://github.com/ergyscokaj/LearningEulersElastica.
@@
Main contributions: This paper presents advancements in the approximation of beam configurations using neural networks. These advancements include: (i) An extensive experimental analysis of approximating numerical discretisations of Euler’s elastica configurations through what we call discrete networks, (ii) Identification and discussion of the limitations associated with this discrete approach, and (iii) Introduction of a new parametrisation strategy called continuous network to address some of these drawbacks.
@@
$\mathcal{L}$
continuous Lagrangian function
$\mathcal{S}$
continuous action functional
$\mathcal{L}_{d}$
discrete Lagrangian function
$\mathcal{S}_{d}$
discrete action functional
$\mathbf{q}$
configuration of the beam
$\mathbf{q}^{\prime}$
first spatial derivative of $\mathbf{q}$
$\theta$
tangential angle
$s$
arc length parameter
$\kappa$
curvature
$L$
length of the undeformed beam
$EI$
bending stiffness, with $E$ the elastic modulus and $I$ the second moment of area
$\hat{\mathbf{q}}$
numerical approximation of $\mathbf{q}$
$N+1$
number of discretisation nodes, with $N$ the number of intervals
$h$
space step (length of each interval)
$q_{\boldsymbol{\rho}}^{\textrm{d}}$
discrete neural network
$q_{\boldsymbol{\rho}}^{\textrm{c}}$
continuous neural network approximating the solution curve $\boldsymbol{q}(s)$
$\theta_{\boldsymbol{\rho}}^{\textrm{c}}$
continuous neural network approximating the angular function $\theta(s)$
$\boldsymbol{\rho}$
parameters of the neural network
$\ell$
number of layers in the neural network
$\sigma$
activation function
$M$
number of training data
$B$
size of one training batch
@@
$\mathcal{D}$
differential operator
$\mathcal{I}$
quadrature operator
@@ 2 Euler’s elastica model
We consider an inextensible beam model in which the cross-section $\Sigma(s)$ is assumed to be constant along the arc length $s$ and perpendicular to the centerline $\mathbf{q}(s)$, which means that no shear deformation can occur. Thus, the deformation of the centerline is a pure bending problem, precisely the Euler’s elastica curve. In the following, we assume $\mathbf{q}\in C^{2}([0,L],\mathbb{R}^{2})$, i.e., the curve is planar and twice continuously differentiable with length $L$. If $s$ denotes the arc length parameter, then $\|\mathbf{q}^{\prime}(s)\|=1$, where ${}^{\prime}=\frac{d}{ds}$, for all $s\in[0,L]$. The elastica problem consists in minimising the following Euler-Bernoulli energy functional
@@
Here, abusing the notation, ${}^{\prime}$ denotes a spatial derivative, but we do not initially assume arc length parametrisation. The parameter $EI$ is the bending stiffness, which governs the response of the elastica under bending. This mechanical parameter consists of a material and a geometric properties, where $E$ is the Young’s modulus and $I$ is the second moment of area of the cross-section $\Sigma$. For simplicity, these parameters are assumed to be constant along the length of the beam.
@@
[…]
@@ 2.1 Space discretisation of the elastica
The continuous augmented Lagrangian $\widetilde{\mathcal{L}}$ in Equation (3) and the action integral $\mathcal{S}$ in Equation (4) are discretised over the beam length $L$ with constant space steps $h$ and $N+1$ equidistant nodes $0=s_{0}<s_{1}<\ldots<s_{N-1}<s_{N}=L$. In second-order systems, the discrete Lagrangian is a function $\widetilde{L}_{d}:TQ\times TQ\times\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. In this study, we refer to a discretisation of the Lagrangian function proposed in [39] based on the trapezoidal rule:
~$\begin{split}&\widetilde{\mathcal{L}}_{d}\left(\mathbf{q}_{k},\mathbf{q}^{\prime}_{k},\mathbf{q}_{k+1},\mathbf{q}^{\prime}_{k+1},\Lambda_{k},\Lambda_{k+1}\right)\\ &=\frac{h}{2}\left[\widetilde{\mathcal{L}}(\mathbf{q}_{k},\mathbf{q}^{\prime}_{k},\left(\mathbf{q}^{\prime\prime}_{k}\right)^{-},\Lambda_{k})+\widetilde{\mathcal{L}}(\mathbf{q}_{k+1},\mathbf{q}^{\prime}_{k+1},\left(\mathbf{q}^{\prime\prime}_{k+1}\right)^{+},\Lambda_{k+1})\right]\end{split}$
where $\mathbf{q}_{k}$, $\mathbf{q}^{\prime}_{k}$ and $\Lambda_{k}$ are approximations of $\mathbf{q}(s_{k})$, $\mathbf{q}^{\prime}(s_{k})$, and $\Lambda(s_{k})$, and the curvature on the interval $[s_{k},s_{k+1}]$ is approximated in terms of lower order derivatives as follows
@@
for $k=1,\dots,N-1$, which approximate the equilibrium equations of the beam in Equations (5) and can be solved together with the boundary conditions.
@@ 2.2 Data generation
Based on these parameters and boundary values, we generate a data set of $2000$ trajectories ($1000$ trajectories for each case) by minimising the particular action in Equation (6), with the trust-constr solver of the optimize.minimize procedure provided in SciPy [42]. We check the resulting solutions by using them as initial guesses for the optimize.root method of SciPy, solving the discrete Euler-Lagrange equations (7).
@@ 3 Approximation with neural networks
We start providing a concise overview on neural networks, and we refer to [43, 25] and references therein for a more comprehensive introduction. A neural network is a parametric function $f_{\boldsymbol{\rho}}:\mathcal{I}\rightarrow\mathcal{O}$ with parameters $\boldsymbol{\rho}\in\Psi$ given as a composition of multiple transformations,
@@
$f_{j}^{MLP}(\mathbf{x})=\sigma(\mathbf{A}_{j}\mathbf{x}+\mathbf{b}_{j})\in\mathbb{R}^{n_{j}},$ (9)
where $\mathbf{x}\in\mathbb{R}^{n_{j-1}}$, and $\mathbf{A}_{j}\in\mathbb{R}^{n_{j}\times n_{j-1}}$, $\mathbf{b}_{j}\in\mathbb{R}^{n_{j}}$ are the parameters of the $j$-th layer, i.e., $\boldsymbol{\rho}=\{\mathbf{A}_{j},\mathbf{b}_{j}\}_{j=1}^{\ell}$. The activation function $\sigma$ is a continuous nonlinear scalar function, which acts component-wise on vectors. The architecture of the neural network is prescribed by the layers $f_{j}$ in Equation (8) and determines the space of functions =$\mathcal{F}=\{f_{\boldsymbol{\rho}}:\mathcal{I}\rightarrow\mathcal{O},\,\,\boldsymbol{\rho}\in\Psi\}$ that can be represented. The weights $\boldsymbol{\rho}$ are chosen such that $f_{\boldsymbol{\rho}}$ approximates accurately enough a map of interest $f:\mathcal{I}\rightarrow\mathcal{O}$. Usually, this choice follows from minimising a purposely designed loss function $\rm{Loss}(\boldsymbol{\rho})$.
@@
Once the training is complete, we assess the model’s accuracy in predicting the correct output for new inputs included in the test set that are unseen during training. In the following, we measure the accuracy on both the training and the test data using the mean squared error of the difference between the predicted trajectories and the true ones.
@@
where $\left(\mathbf{q}^{i}_{0},\mathbf{q}^{i}_{N},(\mathbf{q}^{i}_{0})^{\prime},(\mathbf{q}^{i}_{N})^{\prime}\right)\in\mathbb{R}^{8}$. In order to tackle this problem, we require a set of evaluations $\{\mathbf{q}^{i}_{k},(\mathbf{q}^{i}_{k})^{\prime}\}$ on the nodes $s_{k}\in[0,L]$ of a discretisation. More precisely, in our setting, the data set includes numerical approximations $\hat{\mathbf{q}}$ of the solution $\mathbf{q}(s)$ and its spatial derivative $\mathbf{q}^{\prime}(s)$ at the $N-1$ discrete locations $s_{k}=\frac{kh}{L}$ in the interval $[0,L]$, for $M$ sets of boundary conditions, as described in Section 2.2.

### v2

@@ Abstract
Euler’s elastica is a classical model of flexible slender structures relevant in many industrial applications. Static equilibrium equations can be derived via a variational principle. The accurate approximation of solutions to this problem can be challenging due to nonlinearity and constraints. We here present two neural network-based approaches for simulating Euler’s elastica. Starting from a data set of solutions of the discretised static equilibria, we train the neural networks to produce solutions for unseen boundary conditions. We present a discrete approach learning discrete solutions from the discrete data. We then consider a continuous approach using the same training data set but learning continuous solutions to the problem. We present numerical evidence that the proposed neural networks can effectively approximate configurations of the planar Euler’s elastica for a range of different boundary conditions.
@@ 1 Introduction
Modelling of mechanical systems is relevant in various branches of engineering. Typically, it leads to the formulation of variational problems and differential equations, whose solutions are approximated with numerical techniques. The efficient solution of linear and nonlinear systems resulting from the discretisation of mechanical problems has been a persistent challenge of applied mathematics. While classical solvers are characterised by a well-established and mature body of literature [1, 2, 3, 4, 5, 6, 7], the past decade has witnessed a surge in the use of novel machine learning-assisted techniques [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. These approaches aim at enhancing solution methods by leveraging the wealth of available data and known physical principles. The use of deep learning techniques to improve the performance of traditional numerical algorithms in terms of efficiency, accuracy, and computational scalability [9], is becoming increasingly popular also in computational mechanics (see, e.g. [25]). Examples include a wide range of problems that require the approximation of functions, as well as efficient reduced order modeling [26] or more specific numerical tasks such as optimising the quadrature rule for computing the finite element stiffness matrix [27] or the investigation of data-driven numerical frameworks for the bifurcation analysis of partial differential equations [28, 29]. This recent literature is evidence that neural networks can be used successfully as surrogate models for the solution operators of various differential equations.
@@
In the context of ordinary and partial differential equations, two main trends can be identified. The first one aims at providing a machine learning-based approximation to the discrete solutions of differential problems on a specific space-time grid, for example, by solving linear or nonlinear systems efficiently and accelerating convergence of iterative schemes [20, 19, 15, 14, 13]. The second one provides instead solutions to the differential problem as continuous (and differentiable) functions of the temporal and spatial variables. Depending on the context, conditions on such approximate solutions are provided by the differential problem itself, the initial values and boundary conditions, and the available data. The idea of providing approximate solutions as functions defined on the space-time domain and parametrised as neural networks was proposed in the nineties [30] and was recently revived in the framework of Physics-Informed Neural Networks in [10]. Since then, such an approach has attracted much interest and developed in many directions [8, 21, 31].
@@
In this work, we use neural networks to approximate the configurations of highly flexible slender structures modelled as beams. Such models are of great interest in industrial applications like cable car ropes, diverse types of wires or endoscopes [32, 33, 34, 35]. Notwithstanding their ingenious and simple mathematical formulation, slender structure models can accurately reproduce complex mechanical behaviour and for this reason their numerical discretisation is often challenging. Furthermore, the use of 3-dimensional models requires high computational time. Due to the fact that slender deformable structures have one dimension (length) being orders of magnitude larger than their other dimensions (cross-section), it is possible to reduce the complexity of the problem from a $3$-dimensional elastic continuum to a $1$-dimensional beam. A beam is modelled as a centerline curve, $\mathbf{q}:[0,L]\to\mathbb{R}^{n},s\mapsto\mathbf{q}(s)$, with $n=2$ or $n=3$, along which a rigid cross-section $\Sigma(s)$ is attached. The main model assumption is that the diameter of $\Sigma(s)$ is small compared with the undeformed length $L$. The complexity of the model depends on factors such as the dimension of the problem, the translational and rotational degrees of freedom (DOF) at each node of the beam, the analysis, i.e., static or dynamic. Exploring the numerous beam models documented in the literature, we choose to approach the challenge of approximating beam deformations using a simple yet widely employed model, i.e., the $2$-dimensional Euler’s elastica [36]. The cross-section $\Sigma(s)$ is assumed to have unchanged geometrical and material properties, and be orthogonal to the centerline $\mathbf{q}(s)$. The latter is an inextensible curve and solution of a bending energy minimisation problem [37, 38, 39] for given boundary conditions.
Although the 2-dimensional Euler’s elastica is relatively simple compared to more comprehensive models, it can robustly represent interesting real-world phenomena. For instance, the elastica model appropriately captures the high bending deformations of flexible endoscopes, complex medical devices, during surgeries [33]. The approximation of the elastica through neural networks can help predict the deformed configuration of the beam for endoscopy simulations, particularly when the beam encounters constraints in confined spaces.
@@
When approximating static equilibria of Euler’s elastica via neural networks, a key issue is to ensure the inextensibility of the curve (having unit norm tangents) as well as the boundary conditions. Two main approaches can be found in the literature [31, 21, 40]. One is the weak imposition of constraints and boundary conditions adding appropriate extra terms to the loss function. The other is a strong imposition strategy consisting in shaping the network architectures to satisfy the constraints by construction. We show examples of both the approaches in Sections 4 and 5.
@@
The paper is organised as follows. In Section 2, we present the mathematical model of the planar Euler’s elastica, including its continuous and discrete equilibrium equations. We describe the approach used to generate the data sets for the numerical experiments. In Section 3, we introduce some basic theory and notation for neural networks that we shall use in the succeeding sections. Starting from general theory, we specialise in the task of approximating configurations of Euler’s elastica. In Section 4, we introduce the discrete approach, which aims to approximate precomputed numerical discretisations of Euler’s elastica. This represents the natural approach to approximate the discrete solution trajectories with a parametric method. We discuss some drawbacks associated with this approach and then propose an alternative approximation strategy in Section 5, that leverages the fact that we are approximating a continuous curve on a spatial grid. The continuous approach consists in computing an arc length parametrisation of the beam configuration. We provide insights into two additional networks and analyse how the test accuracy changes with varying constraints, such as boundary conditions or tangent vector norms. Data and codes for the numerical experiments are available in the GitHub repository associated to the paper [^1] https://github.com/ergyscokaj/LearningEulersElastica.
@@
Main contributions: This paper presents advancements in the approximation of beam static configurations using neural networks. These advancements include: (i) A detailed experimental analysis of approximating numerical discretisations of Euler’s elastica configurations through what we call discrete network, (ii) Identification and discussion of the limitations associated with this discrete approach, and (iii) Introduction of a new parametrisation strategy called continuous network to address some of these drawbacks.
@@
$\mathcal{L}$ continuous Lagrangian function
$\mathcal{S}$ continuous action functional
$\mathcal{L}_{d}$ discrete Lagrangian function
$\mathcal{S}_{d}$ discrete action functional
$\mathbf{q}$ configuration of the beam
$\mathbf{q}^{\prime}$ first spatial derivative of $\mathbf{q}$
$\theta$ tangential angle
$s$ arc length parameter
$\kappa$ curvature
$L$ length of the undeformed beam
$EI$ bending stiffness, with $E$ the elastic modulus and $I$ the second moment of area
$\hat{\mathbf{q}}$ numerical approximation of $\mathbf{q}$
$N+1$ number of discretisation nodes, with $N$ the number of intervals
$h$ space step (length of each interval)
$q_{\boldsymbol{\rho}}^{\textrm{d}}$ discrete neural network
$q_{\boldsymbol{\rho}}^{\textrm{c}}$ continuous neural network approximating the solution curve $\boldsymbol{q}(s)$
$\theta_{\boldsymbol{\rho}}^{\textrm{c}}$ continuous neural network approximating the angular function $\theta(s)$
$\boldsymbol{\rho}$ parameters of the neural network
$\ell$ number of layers in the neural network
$\sigma$ activation function
$M$ number of training data
$B$ size of one training batch
@@
$\mathcal{D}$ differential operator
$\mathcal{I}$ quadrature operator
@@ 2 Euler’s elastica model
We consider an inextensible beam model in which the cross-section $\Sigma(s)$ is assumed to be constant along the arc length $s$ and perpendicular to the centerline $\mathbf{q}(s)$, which means that no shear deformation can occur. Thus, the deformation of the centerline is a pure bending problem, precisely Euler’s elastica curve. In the following, we assume $\mathbf{q}\in C^{2}([0,L],\mathbb{R}^{2})$, i.e., the curve is planar and twice continuously differentiable with length $L$. If $s$ denotes the arc length parameter, then $\|\mathbf{q}^{\prime}(s)\|=1$, where ${}^{\prime}=\frac{d}{ds}$, for all $s\in[0,L]$. The elastica problem consists in minimising the following Euler-Bernoulli energy functional
@@
Here, abusing the notation, denotes a spatial derivative, but we do not initially assume arc length parametrisation. The parameter $EI$ is the bending stiffness, which governs the response of the elastica under bending. This mechanical parameter consists of a material and a geometric properties, where $E$ is the Young’s modulus and $I$ is the second moment of area of the cross-section $\Sigma$. For simplicity, these parameters are assumed to be constant along the length of the beam.
@@
which need to be satisfied together with the boundary conditions on positions and tangents, i.e., $(\mathbf{q}(0),\mathbf{q}^{\prime}(0))=(\mathbf{q}_{0},\mathbf{q}^{\prime}_{0})$ and $(\mathbf{q}(L),\mathbf{q}^{\prime}(L))=(\mathbf{q}_{N},\mathbf{q}^{\prime}_{N})$.
@@ 2.1 Space discretisation of the elastica
The continuous augmented Lagrangian $\widetilde{\mathcal{L}}$ in Equation (3) and the action integral $\mathcal{S}$ in Equation (4) are discretised over the beam length $L$ using constant step size $h=L/N$, with $N+1$ the number of the resulting equidistant nodes $0=s_{0}<s_{1}<\ldots<s_{N-1}<s_{N}=L$. In second-order systems, the discrete Lagrangian is a function ~$\widetilde{\mathcal{L}}_{d}:TQ\times TQ\times\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. In this study, we refer to a discretisation of the Lagrangian function proposed in [42] based on the trapezoidal rule:
~$\begin{split}&\widetilde{\mathcal{L}}_{d}\left(\mathbf{q}_{k},\mathbf{q}^{\prime}_{k},\mathbf{q}_{k+1},\mathbf{q}^{\prime}_{k+1},\Lambda_{k},\Lambda_{k+1}\right)\\ &=\frac{h}{2}\left[\widetilde{\mathcal{L}}(\mathbf{q}_{k},\mathbf{q}^{\prime}_{k},\left(\mathbf{q}^{\prime\prime}_{k}\right)^{-},\Lambda_{k})+\widetilde{\mathcal{L}}(\mathbf{q}_{k+1},\mathbf{q}^{\prime}_{k+1},\left(\mathbf{q}^{\prime\prime}_{k+1}\right)^{+},\Lambda_{k+1})\right],\end{split}$
where $\mathbf{q}_{k}$, $\mathbf{q}^{\prime}_{k}$, and $\Lambda_{k}$ are approximations of $\mathbf{q}(s_{k})$, $\mathbf{q}^{\prime}(s_{k})$, and $\Lambda(s_{k})$, and the curvature on the interval $[s_{k},s_{k+1}]$ is approximated in terms of lower order derivatives as follows
@@
for $k=1,\dots,N-1$, which approximate the equilibrium equations of the beam in Equations (5) and can be solved together with the boundary conditions. Here, $D_{i}$ for $i=1,\dots,6$ denotes the differentiation with respect to the $i$-th argument.
@@ 2.2 Data generation
Based on these parameters and boundary values, and using cubic splines as initial guess, we generate a data set of $2000$ trajectories ($1000$ trajectories for each case) by minimising the particular action in Equation (6), with the trust-constr solver of the optimize.minimize procedure provided in SciPy [45]. We check the resulting solutions by using them as initial guesses for the optimize.root method of SciPy, solving the discrete Euler-Lagrange equations (7).
The learning problem we consider relies on numerically generated solution curves. This choice allows us to work with data points that are quantifiably close to the analytical solution of Euler’s elastica. Consequently, showing that the neural networks we propose can accurately approximate these curves translates into their ability to approximate the analytical solution accurately. The motivation of this strategy is not to improve on the numerical solver we use, but to use its accuracy to train a model able to extrapolate to unseen boundary conditions and generate their solution curves more efficiently than the numerical method itself. The chosen supervised learning setting is independent of the fact that we use numerical solutions as data. Indeed, if one had another reliable approximation of the analytical solution, for example, based on realistic measurements, those could also be used or combined with numerically generated trajectories. Using numerical solutions as data is not an inherent limitation of the proposed procedure but a choice we make to quotient out the issue of not having reliable input data. Furthermore, we mainly focus on the development of neural networks able to approximate such input data with high accuracy.
@@ 3 Approximation with neural networks
We start by providing a concise overview of neural networks, which also serves to define the notation used in Sections 4 and 5. We refer to [46, 31, 47] and references therein for a more extensive introduction. A neural network is a parametric function $f_{\boldsymbol{\rho}}:\mathcal{I}\rightarrow\mathcal{O}$ with parameters $\boldsymbol{\rho}\in\Psi$ given as a composition of multiple transformations,
@@
$f_{j}^{\mathrm{MLP}}(\mathbf{x})=\sigma(\mathbf{A}_{j}\mathbf{x}+\mathbf{b}_{j})\in\mathbb{R}^{n_{j}},$ (9)
where $n_{j}$ is the dimension of the output of the $j$-th layer, $\mathbf{x}\in\mathbb{R}^{n_{j-1}}$, and $\mathbf{A}_{j}\in\mathbb{R}^{n_{j}\times n_{j-1}}$, $\mathbf{b}_{j}\in\mathbb{R}^{n_{j}}$ are the parameters of the $j$-th layer, i.e., $\boldsymbol{\rho}=\{\mathbf{A}_{j},\mathbf{b}_{j}\}_{j=1}^{\ell}$. The activation function $\sigma$ is a continuous nonlinear scalar function, which acts component-wise on vectors. The architecture of the neural network is prescribed by the layers $f_{j}$ in Equation (8) and determines the space of functions =$\mathcal{F}=\{f_{\boldsymbol{\rho}}:\mathcal{I}\rightarrow\mathcal{O},\,\,\boldsymbol{\rho}\in\Psi\}$ that can be represented. The weights $\boldsymbol{\rho}$ are chosen such that $f_{\boldsymbol{\rho}}$ approximates accurately enough a map of interest $f:\mathcal{I}\rightarrow\mathcal{O}$. Usually, this choice follows from minimising a purposely designed loss function $\rm{Loss}(\boldsymbol{\rho})$.
@@
During training, we evaluate the model’s prediction accuracy using inputs in a validation set. This helps to prevent overfitting on the training data and may serve as a stopping criterion if the training loss diminishes but the validation error rises. Once the training is complete, we assess the model’s accuracy in predicting the correct output for new inputs included in a test set composed of boundary conditions outside the training and validation sets. In the following, we measure the accuracy on the training, validation, and test data using the mean squared error of the difference between the predicted trajectories and the true ones.
@@
where $\left(\mathbf{q}^{i}_{0},\mathbf{q}^{i}_{N},(\mathbf{q}^{i}_{0})^{\prime},(\mathbf{q}^{i}_{N})^{\prime}\right)\in\mathbb{R}^{8}$. To tackle this problem, we require a set of evaluations $\{\mathbf{q}^{i}_{k},(\mathbf{q}^{i}_{k})^{\prime}\}$ on the nodes $s_{k}\in[0,L]$ of a discretisation. More precisely, in our setting, the data set includes numerical approximations $\hat{\mathbf{q}}$ of the solution $\mathbf{q}(s)$ and its spatial derivative $\mathbf{q}^{\prime}(s)$ at the $N-1$ discrete locations $s_{k}=\frac{kh}{L}$ in the interval $[0,L]$, for $M$ pairs of boundary conditions, as described in Section 2.2.